### v1

# On quasi-Albanese morphisms for log canonical Calabi-Yau pairs

Abstract.
We study the quasi-Albanese morphisms for log canonical Calabi-Yau pairs.
2020 Mathematics Subject Classification:
14E30
1. Introduction
Let $X$ be a complex projective variety with canonical singularities and $K_{X}$ torsion. Kawamata proved that its Albanese morphism $\alpha:X\to A$ is an étale fiber bundle [Kaw85, Theorem 8.3]. Inspired by the works [Iit76, Kaw81, Zha05, Fuj24, BFPT24], we study quasi-Albanese morphisms for (the open part of) log canonical Calabi-Yau pairs.
Before stating the main theorem, we recall the definition of logarithmic Kodaira dimension and a theorem by Kawamata and Fujino. Let $X^{0}$ be a smooth quasi-projective variety. Let $X$ be a compactification of $X^{0}$ such that $D_{X}:=X\setminus X^{0}$ is a simple normal crossing divisor. The logarithmic Kodaira dimension $\overline{\kappa}(X^{0})$ of $X^{0}$ is defined as $\kappa(K_{X}+D_{X})$. By Iitaka [Iit77], $\overline{\kappa}(X^{0})$ is independent of the choice of such $X$.
Theorem 1.
[Kaw81, Thoerem 28], see also [Fuj24, Theorem 1.2] If $\overline{\kappa}(X^{0})=0$, then the quasi-Albanese morphism of $X^{0}$ is dominant and has irreducible general fibers.
The following is our main theorem.
Theorem 2.
Let $(X,D_{X})$ be a log smooth pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $X^{0}=X\setminus D_{X}$ and let $\alpha:X^{0}\to G$ be its quasi-Albanese morphism. Then
  1. (1)
    $\alpha$ is surjective in codimension one: for any prime divisor $P\subset G$, there is a prime divisor $Q\subset X^{0}$ dominating $P$,
  2. (2)
    $\alpha$ is flat in codimension one: for any prime divisor $Q\subset X^{0}$, if $Q$ do not dominate $G$, then $\alpha(Q)$ is of codimension one,
  3. (3)
    $\alpha$ is semistable in codimension one: for any prime divisor $P\subset G$, the pair $(X,\alpha^{*}P)$ is log canonical near the generic point of $P$.
Theorem 2 is a consequence of the following theorem, whose proof depends on Theorem 1 and the techniques from the minimal model program.
Theorem 3.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Set $q(X):=\dim H^{1}(O_{X})$ and $q(X,D_{X}):=\dim H^{0}(\Omega_{X^{0}}^{1}(\log D_{X}|_{X^{0}}))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$. Let $A$ be the Albanese variety of $X$ and let $G$ be the quasi-Albanese variety of $X\setminus D_{X}$. Then there are
  • a $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$,
  • a crepant birational contraction $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$,
  • $(q(X,D_{X})-q(X))$’s line bundles $L_{i}\in\mathrm{Pic}^{0}(A)$, and
  • a commutative diagram,
[…]
    where $\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$ is a compactification of $G$, $H=\mathbb{P}_{A}\setminus G$ is a simple normal crossing divisor, $g$ is a fibration, and $p$ is the natural projection,
such that the following holds.
  1. (1)
    $p\circ g$ is the Albanese morphism of $W$, $h$ is the quasi-Albanese morphism of $W\setminus D_{W}$.
  2. (2)
    The $g$-vertical part of $D_{W}$ coincides with $g^{-1}H$.
  3. (3)
    The canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ has the form
    $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
    and the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable.
As a corollary, if $q(X)=0$ and $q(X,D_{X})=\dim X$, then $(X,D_{X})$ is crepant birational to $(\mathbb{P}^{n},(\prod_{i}z_{i}=0))$. Another corollary is that if $q(X,D_{X})-q(X)\geq 2$, then $D_{X}$ is connected.
In the previous Theorem, the pair $(W,D_{W})$ is closely related to $(X,D_{X})$. For example, they have the same orbifold fundamental group, and the same dual complex (up to a PL-homeomorphism) (cf. [dFKX17, Proposition 11]). On the other hand, the fibration $g:(W,D_{W})\to\mathbb{P}_{A}$ has nice properties, which may reduce some questions about $(X,D_{X})$ to the case $q(X,D_{X})=0$ by induction on dimension. We give an example in the last section.
Acknowledgments
The author thanks professors Zhan Li, Lei Zhang, and Junchao Shentu for useful discussions. The author is supported by USTC.
2. Preparations
We work over the field of complex numbers. For the singularities of pairs and the minimal model program, we refer to the book [KM98].
A projective surjective morphism between normal varieties with connected fibers is called a fibration. Let $f:X\to Y$ be a fibration between normal varieties. Let $D$ be an effective $\mathbb{Q}$-divisor on $X$. We can write $D=D^{h}+D^{v}$, where $D^{v}$ does not dominate $Y$, and each component of $D^{h}$ dominates $Y$. $D^{v}$ (resp. $D^{h}$) is called the vertical part (resp. horizontal) part of $D$. $D$ is called vertical (resp. horizontal) if $D=D^{v}$ (resp. $D=D^{h}$). Assume that $D$ is vertical, then $D$ is called exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and very exceptional over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $Q$ dominates $P$ but $Q\not\subset D$.
For a birational map $\pi:X\dashrightarrow Y$ between normal varieties, we set $\operatorname{Ex}(\pi):=\sum P$, where the sum is taken over all prime divisors on $X$ which are exceptional over $Y$. A birational map $\pi:X\dashrightarrow Y$ is called a birational contraction if $\operatorname{Ex}(\pi^{-1})=0$.
We will use the following fact implicitly: Let $A$, $B$ be two effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisors on a projective variety $X$. If $\operatorname{Supp}A=\operatorname{Supp}B$, then $\kappa(A)=\kappa(B)$. If $\operatorname{Supp}A\subset\operatorname{Supp}B$, then $\kappa(A)\leq\kappa(B)$. Let $f:Y\to X$ be a projective surjective morphism from another projective variety, then $f^{-1}\operatorname{Supp}A=\operatorname{Supp}f^{*}A$.
For a projective variety $X$, we set $q(X):=\dim H^{1}(O_{X})$, and for a projective lc pair $(X,D_{X})$, we set $q(X,D_{X}):=\dim H^{0}(\Omega_{X^{0}}^{1}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$.
2.1. Differential forms on lc pairs
Theorem 4 ([GKKP11]).
Let $(X,D_{X})$ be an lc pair. Let $X^{0}$ be its log smooth locus and let $\mathrm{nonklt}(X,D_{X})$ be its non-klt locus. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}$ be the codimension one part of $\pi^{-1}\mathrm{nonklt}(X,D_{X})$. Then $\pi_{*}\Omega_{Y}^{k}(\log D_{Y})$ is reflexive and $H^{0}(Y,\Omega_{Y}^{k}(\log D_{Y}))\cong H^{0}(X^{0},\Omega_{X^{0}}^{k}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$ for all $k$.
Proposition 5.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then
  1. (1)
    $q(X)=q(Y)$,
  2. (2)
    $q(X,D_{X})=q(Y,D_{Y})$,
  3. (3)
    $\kappa(K_{X}+D)=\kappa(K_{Y}+D_{Y})$ and $\nu(K_{X}+D)=\nu(K_{Y}+D_{Y})$,
where $\kappa$ and $\nu$ are Kodaira dimension and numerical Kodaira dimension respectively.
Proof.
Since klt varieties have rational singularities, we have (1). Since $(X,D_{X})$ is lc, there is an effective exceptional divisor $E$ on $Y$ such that $K_{Y}+D_{Y}=\pi^{*}(K_{X}+D_{X})+E$, so we have (3).
Now we prove (2). Note that $D_{Y}\geq\pi^{-1}D_{X}$ and $D_{X}=\mathrm{nonklt}(X,D_{X})$. By Theorem 4, to show (2), it suffices to show that the injective homomorphism
$\iota:H^{0}(\Omega_{Y}^{1}(\log\pi^{-1}D_{X}))\to H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$
is an isomorphism. Write $\pi^{-1}D_{X}=\sum_{i}P_{i}$ and $D_{Y}-\pi^{-1}D_{X}=\sum_{j}E_{j}$ as sum of prime divisors. By the exact sequence
$0\to\Omega_{Y}^{1}\to\Omega_{Y}^{1}(\log(\sum_{i}P_{i}+\sum_{j}E_{j}))\to\oplus_{i}O_{P_{j}}\oplus\oplus_{j}O_{E_{j}}\to 0,$
one checks that $\iota$ is an isomorphism iff $\sum_{i}a_{i}c_{1}(O_{Y}(P_{i}))+\sum_{j}b_{j}c_{1}(O_{Y}(E_{j}))=0$ implies that all $b_{j}$’s are zero. Which is indeed the case since all $E_{j}$’s are exceptional and $X$ is $\mathbb{Q}$-factorial.∎
Lemma 6.
Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective $\mathbb{Q}$-factorial dlt pairs with reduced $D_{X}$ and $D_{Y}$. Assume that there is a birational contraction $\pi:X\dashrightarrow Y$ such that $D_{X}=\pi^{-1}_{*}D_{Y}+\operatorname{Ex}(\pi)$ then $q(X,D_{X})=q(Y,D_{Y})$.
Proof.
Let $X\xleftarrow{p}W\xrightarrow{q}Y$ be a common log resolution. Then $q^{-1}_{*}D_{Y}+\operatorname{Ex}(q)=p^{-1}_{*}D_{X}+\operatorname{Ex}(p)$. Hence, the Lemma follows from the previous proposition. ∎
2.2. Quasi-Albanese morphism
Theorem 7 (Quasi-Albanese morphism [Iit76], see also [Fuj24]).
Let $X$ be a smooth quasi-projective variety. Then there is a morphism $\alpha:X\to G$ to a quasi-abelian variety such that
(1) for any morphism $\alpha^{\prime}:X\to G^{\prime}$ to a quasi-abelian variety, there is a morphism $h:G\to G^{\prime}$ such that $\alpha^{\prime}=h\circ\alpha$, and
(2) $\alpha$ is uniquely determined.
8Construction.
We recall the constructions of quasi-Albanese morphisms in [Iit76, Fuj24]. Let $(X,D_{X}=\sum_{i=1}^{m}D_{i})$ be a projective log smooth pair with $D_{X}$ reduced. Where $D_{i}$’s are prime divisors. Let $X^{0}:=X\setminus D_{X}$ and let $\iota:X^{0}\to X$ be the open embedding. Let $p:X\to A$ be the Albanese morphism. Set $q=q(X)$, $\overline{q}=q(X,D_{X})$, and $d=\overline{q}-q$.
Let $\eta_{i}$, $1\leq i\leq d$, be a basis of the free part of $\ker(\iota_{*}:H_{1}(X^{0},\mathbb{Z})\to H_{1}(X,\mathbb{Z}))$, and let $\xi_{i}\in H_{1}(X^{0},\mathbb{Z})$, $1\leq i\leq 2q$, be such that these $\eta_{i}$’s and $\xi_{i}$’s form a basis of the free part of $H_{1}(X^{0},\mathbb{Z})$.
Let $\omega_{i}$, $1\leq i\leq q$, be a basis of $H^{0}(\Omega_{X}^{1})$, and let $\phi_{i}\in H^{0}(\Omega_{X}^{1}(\log D))$, $1\leq i\leq d$, be such that these $\omega_{i}$’s and $\phi_{i}$’s form a basis of $H^{0}(\Omega_{X}^{1}(\log D_{X}))$, and satisfy $\int_{\eta_{i}}\omega_{j}=0$ and $\int_{\eta_{i}}\phi_{j}=\delta_{ij}$ (the Kronecker delta).
Set
$A_{i}:=(\int_{\xi_{i}}\omega_{1},\cdots,\int_{\xi_{i}}\omega_{q},\int_{\xi_{i}}\phi_{1},\cdots,\int_{\xi_{i}}\phi_{d})=:(\hat{A}_{i},C_{i}),1\leq i\leq 2q,\text{ and }$
$B_{i}:=(\int_{\eta_{i}}\omega_{1},\cdots,\int_{\eta_{i}}\omega_{q},\int_{\eta_{i}}\phi_{1},\cdots,\int_{\eta_{i}}\phi_{d})=:(0,\hat{B}_{i}),1\leq i\leq d.$
The quasi-Albanese variety of $X^{0}$ is given by the quasi-abelian variety
$G:=\mathbb{C}^{\overline{q}}/\sum_{i}\mathbb{Z}A_{i}+\sum_{j}\mathbb{Z}B_{j}.$
We have the following Chevalley decomposition of $G$:
$0\to\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}\to G\to\mathbb{C}^{q}/\sum_{i}\mathbb{Z}\hat{A}_{i}=A\to 0,$
and $A$ coincides with the Albanese variety of $X$. The isomorphism $\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}$ is given by
$(x_{1},\cdots,x_{d})\mapsto(\exp(2\pi\mathrm{i}x_{1}),\cdots,(\exp(2\pi\mathrm{i}x_{d})).$
Fix a point $x_{0}\in X^{0}$. The quasi-Albanese morphism of $X^{0}$ is then given by
$X^{0}\to G,x\mapsto\left(\int_{x_{0}}^{x}\omega_{1},\cdots,\int_{x_{0}}^{x}\omega_{q},\int_{x_{0}}^{x}\phi_{1},\cdots,\int_{x_{0}}^{x}\phi_{d}\right).$
Remark 9.
Keep the notations introduced in the previous construction. Let $G^{\prime}\to\mathbb{C}^{q}$ be the pullback of the $(\mathbb{C}^{*})^{d}$-bundle $G\to A$ by the cover $\mathbb{C}^{q}\to A$. Then we have $G^{\prime}=\mathbb{C}^{\overline{q}}/\sum_{i}\mathbb{Z}B_{i}\cong\mathbb{C}^{q}\times(\mathbb{C}^{*})^{d}$ and
$G=\mathbb{C}^{q}\times(\mathbb{C}^{*})^{d}/\pi_{1}(A),$
where $\pi_{1}(A)=\sum_{i}\mathbb{Z}{\hat{A}_{i}}$ acts on $\mathbb{C}^{q}\times(\mathbb{C}^{*})^{d}$ diagonally by the group homomorphism:
$\pi_{1}(A)\to\operatorname{Aut}((\mathbb{C}^{*})^{d}),\hat{A}_{i}\mapsto\left((t_{1},\cdots,t_{d})\mapsto(t_{1}\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{1}),\cdots,t_{d}\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{d}))\right),1\leq i\leq q.$
For each $j$, let $L_{j}\in\mathrm{Pic}^{0}(A)$ be the line bundle corresponding to the representation
(1) $\pi_{1}(A)\to\mathbb{C}^{*},\hat{A}_{i}\mapsto\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{j}),1\leq i\leq q.$
Then $G$ is the fiber product over $A$ of the total spaces of $L_{j}$’s with zero sections removed, in other words,
$G=\operatorname{Spec}_{A}\bigoplus_{(m_{1},\cdots,m_{d})\in\mathbb{Z}^{d}}L_{1}^{-m_{1}}\otimes\cdots\otimes L_{d}^{-m_{d}}.$
$G$ has the following natural compactification:
(2) $G\subset\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1}).$
Let $a_{ij}=2\pi\mathrm{i}\mathrm{Res}_{D_{i}}\phi_{j}$, $1\leq i\leq m$, $1\leq j\leq d$. For each $j$, we have the following.
  1. (1)
    $a_{ij}\in\mathbb{Z}$ for all $i$.
  2. (2)
    $p^{*}L_{j}$ is given by the representation $\pi_{1}(X)\to\mathbb{C}^{*},\gamma\mapsto\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$.
  3. (3)
    $t_{j}(x):=\exp(2\pi\mathrm{i}\int_{x_{0}}^{x}\phi_{j})$ is a rational section of $p^{*}L_{j}$ with $\mathrm{div}(t_{j})=\sum_{i}a_{ij}D_{i}$. In particular, $p^{*}L_{j}\cong O_{X}(\sum_{i}a_{ij}D_{i})$.
Proof.
For each $i$, let $\delta_{i}$ be a small loop around $D_{i}$.
(1) Since $\delta_{i}\in\ker(H_{1}(X^{0},\mathbb{Z})\to H_{1}(X,\mathbb{Z}))$, we have $\delta_{i}=\sum_{k}n_{k}\eta_{k}+t$ for some integers $n_{k}$ and a torsion element $t$. By the residue theorem, we have
$2\pi\mathrm{i}\mathrm{Res}_{D_{i}}\phi_{j}=\int_{\delta_{i}}\phi_{j}=\sum_{k}n_{k}\int_{\eta_{k}}\phi_{j}+\int_{t}\phi_{j}=\sum_{k}n_{k}\delta_{kj}=n_{j}\in\mathbb{Z}.$
(2) By the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$, we mean $\exp(2\pi\mathrm{i}\int_{\gamma^{\prime}}\phi_{j})$, where $\gamma^{\prime}$ is any other loop in $X^{0}$ which is homotopic to $\gamma$. Since $\pi_{1}(X^{0})\to\pi_{1}(X)$ is surjective with kernel generated by $\delta_{i}$’s, $\exp(2\pi\mathrm{i}\int_{\delta_{i}}\phi_{j})=1$ for all $i$, and $\phi_{j}$ is $d$-closed, the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$ is well-defined. Hence, we have (ii) by the representation (1).
(3) By $t_{j}(\gamma x)=\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})t_{j}(x)$ for all $\gamma\in\pi_{1}(X)$, we have $t_{j}$ is a holomorphic nowhere vanishing section of $p^{*}L_{j}$ over $X^{0}$. Let $\{x_{i}\}$ be a coordinate system on an open subset $U$ of $X$ such that $D=(x_{1}x_{2}\cdots x_{c}=0)$. Then
$\phi_{j}|_{U}=\sum_{i=1}^{c}\frac{a_{ij}}{2\pi\mathrm{i}}\frac{dx_{i}}{x_{i}}+\text{some holomorphic function},\text{ and }$
$t_{j}|_{U}=\prod_{i=1}^{c}x_{i}^{a_{ij}}\times\text{some holomorphic function without zeros}.$
Hence $t_{j}$ is a meromorphic, and so rational (by Serre’s GAGA) section of $p^{*}L_{j}$ and $\mathrm{div}(t_{j})=\sum_{i}a_{ij}D_{i}$. ∎
Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of the $(\mathbb{P}^{1})^{d}$-bundle $\mathbb{P}_{A}\to A$, the rational map $X\dashrightarrow\mathbb{P}_{A}$ over $A$ is now given by
$x\mapsto(z_{1},\cdots,z_{d})=(t_{1},\cdots,t_{d})\in\mathbb{P}_{A}.$
Let $H:=\mathbb{P}_{A}\setminus G=\sum_{i}((z_{i}=0)+(z_{i}=\infty))$. The pair $(\mathbb{P}_{A},H)$ is log smooth and $K_{\mathbb{P}_{A}}+H\sim 0$.
Example 10.
Let $(X,D_{X})$ be a log smooth pair with $D_{X}$ reduced.
(1) Let $X=\mathbb{P}^{2}$. If $D_{X}=$ 3 lines, then $q(X,D_{X})=3$. If $D_{X}=$ an elliptic curve, then $q(X,D_{X})=0$. If $D_{X}=Q+L$, where $L=(l=0)$ is a line and $Q=(q=0)$ is a conic, then $q(X,D_{X})=1$, $H^{0}(\Omega^{1}_{X}(\log D_{X}))=\mathbb{C}\cdot d\log\frac{l^{2}}{q}$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},x\mapsto\frac{l^{2}}{q}$.
(2) Let $X=\mathbb{P}^{1}_{z_{1}}\times\mathbb{P}^{1}_{z_{2}}$ and $D_{X}=F_{1}+F_{2}+\Delta$ where $F_{1}=(z_{1}=0)$, $F_{2}=(z_{2}=0)$, $\Delta=(z_{1}=z_{2})$. Then $H^{0}(\Omega_{X}(\log D_{X}))=\mathbb{C}\cdot d\log(1/z_{1}-1/z_{2})$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},(z_{1},z_{2})\mapsto 1/z_{1}-1/z_{2}$.
(3) Let $A$ be an abelian variety and let $L$ be a line bundle on $A$. Let $X=\mathbb{P}_{A}(O_{A}\oplus L)$, let $z$ be its inhomogeneous fiber coordinate, and let $H_{0}=(z=0)$, $H_{\infty}=(z=\infty)$, $D_{X}=H_{0}+H_{\infty}$. Then $q(X)=\dim A$.
If $L\in\text{Pic}^{0}(A)$, then $q(X,D_{X})=\dim A+1$, and $\mathbb{P}_{A}\setminus H_{1}+H_{2}$ itself is a quasi-abelian variety.
If $L\not\in\text{Pic}^{0}(A)$, then $q(X,D_{X})=\dim A$, and the quasi-Albanese morphism is the natural projection $\mathbb{P}_{A}\setminus H_{1}+H_{2}\to A$.
Remark 11.
Let $X$ be a (possibly) singular variety. Let $\pi:X^{\prime}\to X$ be a resolution and let $f^{\prime}:X^{\prime}\to G$ be the quasi-Albanese morphism of $X^{\prime}$, then the quasi-Albanese map of $X$ is defined as the rational map $f^{\prime}\circ\pi^{-1}$. Assume that $X$ has klt singularities. Then the fibers of $\pi$ are rationally chain connected by Hacon-McKernan [HM07]. Since $G$ contains no rational curves, $f^{\prime}$ factors through $\pi$ by the rigidity Lemma (cf.[Deb01, Lemma 1.15]). The statement in [Deb01, Lemma 1.15] requires $f^{\prime}$ to be proper, but the proof only used that the image of $X^{\prime}\to X\times G$ is proper over $X$, which is guaranteed by [Har77, Chapter II Exercise 4.4].
Proposition 12.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $\kappa(K_{X}+D_{X})=0$. Then the quasi-Albanese morphism $X\setminus D_{X}\to G$ is dominant and has irreducible general fibers.
Proof.
Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}:=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then $q(Y,D_{Y})=q(Y,\pi^{-1}D_{X})=q(X,D_{X})$ and $\kappa(K_{Y}+D_{Y})=0$ by Proposition 5. By the universal property of quasi-Albanese morphisms, we have the following commutative diagram.
[…]
where $g$, $g^{\prime}$ and $f$ are quasi-Albanese morphisms. Since the injective homomorphism $H^{0}(\Omega^{1}_{Y}(\log\pi^{-1}D_{X}))\to H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$ is an isomorphism, we have $H^{1}(Y\setminus D_{Y},\mathbb{C})=H^{1}(O_{Y})\oplus H^{0}(\Omega_{Y}(\log\pi^{-1}D_{X}))\to H^{1}(Y\setminus\pi^{-1}D_{X},\mathbb{C})=H^{1}(O_{Y})\oplus H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$ is an isomorphism by Deligne’s theory of mixed Hodge structures. Hence the surjective morphism $H_{1}(Y\setminus D_{Y},\mathbb{Z})/{\mathrm{torsion}}\to H_{1}(Y\setminus\pi^{-1}D_{X},\mathbb{Z})/{\mathrm{torsion}}$ is also an isomorphism, so $G^{\prime}\to G$ is an isomorphism.
Since $\kappa(K_{Y}+D_{Y})=0$, we have $g^{\prime}$ is dominant and has irreducible general fibers by Theorem 1. Hence, $g$, $f$ are dominant and have irreducible general fibers. ∎
2.3. Subadditivity and positivity
Theorem 13.
Let $f:X\to Y$ be a fibration between smooth projective varieties. Let $D_{X}$ and $D_{Y}$ be reduced snc divisors on $X$ and $Y$ respectively. Denote by $F$ a general fiber of $f$. Assume that $f^{-1}D_{Y}\subset D_{X}$. Then the following holds.
(1)[Mae86, Corollary 2] Let $B$ be a big $\mathbb{Q}$-divisor on $Y$, then
$\kappa(K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})+f^{*}B)\geq\kappa(K_{F}+{D_{X}}|_{F})+\dim Y.$
(2) If $\kappa(K_{F}+{D_{X}}|_{F})\geq 0$, then $K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})$ is pseudoeffective.
[…]
Since the $\mathbb{Q}$-divisor $B$ in (1) can be sufficiently small, (2) follows.
Let $X^{0}$ be a smooth quasi-projective variety and let $X^{0}\subset X$ be a compactification such that $D_{X}:=X\setminus X^{0}$ is a simple normal crossing divisor. By the same reasoning as in the definition of $\overline{\kappa}(X^{0})$ (cf. [Fuj24, Lemma 2.3]), we can define the numerical logarithmic Kodaira dimension as $\overline{\nu}(X^{0}):=\nu(K_{X}+D_{X})$. By Fujino’s subadditivity 13(3), we have the following.
Lemma 14.
Let $f:X\to Y$ be a dominant morphism with irreducible general fibers between smooth quasi-projective varieties. Let $F$ be a general fiber of $f$. Then $\overline{\nu}(X)\geq\overline{\nu}(F)+\overline{\kappa}(Y)$.
3. Proof of the main theorems
Proof of Theorem 2.
Let $\alpha:X\setminus D_{X}\to G$ be the quasi-Albanese morphism. Let $G\subset\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$ be the compactification described in remark 9, where $L_{i}$’s are elements in $\text{Pic}^{0}(A)$ and let $p:\mathbb{P}_{A}\to A$ be the natural projection. Then $H:=\mathbb{P}_{A}\setminus G$ is a snc divisor and $K_{\mathbb{P}_{A}}+H\sim 0$.
Let $\pi:Z\to X$ be a log resolution of $(X,D_{X})$ resolving the indeterminacies of the rational map $X\dashrightarrow\mathbb{P}_{A}$ and let $D_{Z}:=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. By construction, we have
(3) $f^{-1}H\subset D_{Z}.$
Write $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$. Since $(X,D_{X})$ is lc, $E_{Z}$ is effective and exceptional over $X$. Then by Theorem 15, $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$, the canonical bundle formula for $(W,D_{W})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H)$, the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and $g^{-1}H=D_{W}^{v}$.
[…]
Since the birational map $Z\dashrightarrow W$ contracts precisely the divisor $E_{Z}\subset\mathrm{Ex}(\pi)$, we have $\mu:W\dashrightarrow X$ is a birational contraction and $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$. Let $X\xleftarrow{u}V\xrightarrow{v}W$ be a common log resolution of $(X,D_{X})$ and $(W,D_{W})$. Then $u^{*}(K_{X}+D_{X})=v^{*}(K_{W}+D_{W})$ by the Negativity Lemma, i.e., $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is crepant. By Lemma 6, we have $q(W,D_{W})=q(X,D_{X})$ and $q(W)=q(X)$. By the universal property of quasi-Albanese morphisms and Proposition 12, the composite morphism $W\setminus D_{W}\to W\setminus g^{-1}H\to G$ is the quasi-Albanese morphism of $W\setminus D_{W}$. By the same reasoning, $p\circ g:W\to A$ is the Albanese morphism of $W$. ∎
Theorem 15.
Let $(Z,D_{Z})$ be a projective log smooth pair with $D_{Z}$ reduced and $\kappa(K_{Z}+D_{Z})=0$. Let $A$ be the Albanese variety of $Z$. Let $G$ be a quasi-abelian variety whose abelian part is $A$. Let
$\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d})$
be a compactification of $G$ where $L_{i}$’s are elements in $\text{Pic}^{0}(A)$. Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of $\mathbb{P}_{A}$. For each $i$, let $p_{i}:\mathbb{P}_{A}\to\mathbb{P}_{A}(O_{A}\oplus L_{i})$ be the projection, and let
$H_{i}=(z_{i}=0)+(z_{i}=\infty)\subset\mathbb{P}_{A}(O_{A}\oplus L_{d})$
be the effective and ample over $A$ divisor. Let $H=\sum_{i}p_{i}^{*}H_{i}$.
Assume that there are
  • a fibration $f:Z\to\mathbb{P}_{A}$ with $f^{-1}H\subset D_{Z}$,
  • a projective $\mathbb{Q}$-factorial dlt pair $(X,D_{X})$ with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, and
  • a birational morphism $\pi:Z\to X$ such that $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$ and
    $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$
    for some effective and exceptional over $X$ divisor $E_{Z}$.
Then the following holds.
  1. (1)
    Any exceptional over $\mathbb{P}_{A}$ divisor is contained in $E_{Z}+f^{-1}H$.
  2. (2)
    Let $P\subset\mathbb{P}_{A}$ be a prime divisor with $P\not\subset H$, then there is a prime divisor $Q\subset Z$ with $Q\not\subset E_{Z}$ dominating $P$.
  3. (3)
    Assume that for any prime divisor $P\subset H$, there is a prime divisor $Q\not\subset E_{Z}$ dominating $P$, then $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$.
  4. (4)
    Assume that $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$. Let $g:W\to\mathbb{P}_{A}$ be the induced fibration. Then the canonical bundle formula for $(W,D_{W})\to\mathbb{P}_{A}$ has the form
    $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
    the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and the vertical part $D_{W}^{v}$ of $D_{W}$ coincides with $g^{-1}H$.
  5. (5)
    $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$.
Proof.
Note that by assumption, $G=\mathbb{P}_{A}\setminus H$, $(\mathbb{P}_{A},H)$ is log smooth, and $K_{\mathbb{P}_{A}}+H\sim 0$. Let $F$ be a general fiber of $f$. By Fujino’s subadditivity Theorem 13(3), we have
$0=\nu(K_{Z}+D_{Z})\geq\nu(K_{F}+D_{Z}|_{F})+\kappa(K_{\mathbb{P}_{A}}+H),$
so $\nu(K_{F}+D_{Z}|_{F})=0$. By Gongyo’s abundance theorem [Gon11], $(F,D_{Z}|_{F})$ has a good minimal model and $\kappa(K_{F}+D_{Z}|_{F})=0$.
(1) Let $Q\subset Z$ be an exceptional over $\mathbb{P}_{A}$ prime divisor which is not contained in $f^{-1}H$. Let $q:\mathbb{P}^{\prime}_{A}\to\mathbb{P}_{A}$ be a composition of the blowup along $f(Q)$ with a log resolution. We may assume that $q$ is an isomorphism over $\mathbb{P}_{A}\setminus f(Q)$. Let $p:Z^{\prime}\to Z$ be a log resolution of $(Z,D_{Z})$ resolving the indeterminacies of $Z\dashrightarrow\mathbb{P}_{A}^{\prime}$. We have the following commutative diagram.
[…]
Write $K_{Z^{\prime}}+D_{Z^{\prime}}=p^{*}(K_{Z}+D_{Z})+E_{1}$, $K_{\mathbb{P}^{\prime}_{A}}+H^{\prime}=q^{*}(K_{\mathbb{P}_{A}}+H)+E_{2}$, where $H^{\prime}=q^{-1}H$, $D_{Z^{\prime}}=p^{-1}_{*}D_{Z}+\operatorname{Ex}(p)$, and $E_{1}$, $E_{2}$ are effective and exceptional divisors. By $f^{\prime-1}H^{\prime}\subset D_{Z^{\prime}}$,
$0\sim_{\mathbb{Q}}p^{*}(K_{Z}+D_{Z}-f^{*}(K_{\mathbb{P}_{A}}+H)-E_{Z})=K_{Z^{\prime}}+D_{Z^{\prime}}-f^{\prime*}(K_{\mathbb{P}^{\prime}_{A}}+H^{\prime})-E_{1}+f^{\prime*}E_{2}-p^{*}E_{Z},$
and Theorem 13(1), we have $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is pseudoeffective. Since $(\pi\circ p)_{*}(E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z})=-(\pi\circ p)_{*}f^{\prime*}E_{2}$ is pseudoeffective, we have $f^{\prime*}E_{2}$ is exceptional over $X$. Hence $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is effective by a Lemma due to Lazarsfeld (cf. [KL09, Corollary 13]).
Let $Q^{\prime}$ be the strict transform of $Q$ on $Z^{\prime}$ and let $P^{\prime}=f^{\prime}(Q^{\prime})$. Since $q(P^{\prime})=f(Q)\not\subset H$, we have $\operatorname{mult}_{P^{\prime}}E_{2}>0$, so $\operatorname{mult}_{Q^{\prime}}f^{\prime*}E_{2}>0$. By
$0\leq\operatorname{mult}_{Q^{\prime}}(E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z})=\operatorname{mult}_{Q^{\prime}}(-f^{\prime*}E_{2}+p^{*}E_{Z}),$
we have $Q^{\prime}\subset p^{*}E_{Z}$. Hence $Q\subset E_{Z}$.
(2) Assume the contrary, i.e., assume that
$f^{-1}P\subset E_{Z}+E$
for some effective exceptional over $\mathbb{P}_{A}$ divisor $E$. Then by (1), we have
$f^{-1}(P+H)\subset E_{Z}+f^{-1}H\subset D_{Z}.$
Then by Lemma 14 and Lemma 16, we have
$0=\nu(K_{Z}+D_{Z})=\overline{\nu}(Z\setminus D_{Z})\geq\overline{\nu}(F\setminus D_{Z})+\overline{\kappa}({\mathbb{P}_{A}}\setminus P+H)=\overline{\kappa}(G\setminus P)>0,$
which is a contradiction.
(3) By (2) and the assumption in (3), we known that the vertical part $E_{Z}^{v}$ of $E_{Z}$ is very exceptional over $\mathbb{P}_{A}$. Since $\kappa(E_{Z}|_{F})=\kappa(K_{F}+D_{Z}|_{F})=0$, we have
(4) $(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$
for all sufficiently divisible $m>0$, where $\wedge$ means taking reflexive hull.
By Gongyo’s abundance theorem and Hacon-Xu [HX13, Theorem 2.12], $(Z,D_{Z})$ has a good minimal model over a nonempty Zariski open subset $U$ of $\mathbb{P}_{A}$. We run a $(Z,D_{Z})$-MMP over $\mathbb{P}_{A}$ with scaling of an ample divisor. Then this MMP will terminate over $U$ by [HX13, Corollary 2.9]. Hence we have a partial MMP $(Z,D_{Z})\dashrightarrow(Z^{\prime},D_{Z^{\prime}})$ with $K_{Z^{\prime}}+D_{Z^{\prime}}$ semiample over $U$, where $D_{Z^{\prime}}$ is the pushforward of $D_{Z}$. Let $F^{\prime}$ be a general fiber of $f^{\prime}:Z^{\prime}\to\mathbb{P}_{A}$ and let $E_{Z^{\prime}}$ be the pushforward of $E_{Z}$. Since $\kappa(K_{F^{\prime}}+D_{Z^{\prime}}|_{F^{\prime}})=0$, we have $E_{Z^{\prime}}|_{F^{\prime}}\sim_{\mathbb{Q}}K_{Z^{\prime}}+D_{Z^{\prime}}|_{F^{\prime}}\sim_{\mathbb{Q}}0$, so $E_{Z^{\prime}}$ is vertical over $\mathbb{P}_{A}$. By formula (4), we have $(f^{\prime}_{*}O_{Z^{\prime}}(mE_{Z^{\prime}}))^{\wedge}=(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$ for all sufficiently divisible $m>0$. Hence $E_{Z^{\prime}}$ is very exceptional very $\mathbb{P}_{A}$ by [GL13, Lemma 2.15]. Then $(Z^{\prime},D_{Z^{\prime}})$, and so $(Z,D_{Z})$, has a good minimal model over $\mathbb{P}_{A}$ by Birkar [Bir12, Theorem 1.8].
(4) Write the canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ as $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+B_{\mathbb{P}_{A}}+M_{\mathbb{P}_{A}})$. Since $f^{-1}H\subset D_{Z}$, we have $g^{-1}H\subset D_{W}$, so $B_{\mathbb{P}_{A}}\geq H$. Since $K_{W}+D_{W}\sim_{\mathbb{Q}}0$ and $K_{\mathbb{P}_{A}}+H\sim 0$, we have $B_{\mathbb{P}_{A}}=H$ and $M_{\mathbb{P}_{A}}\sim_{\mathbb{Q}}0$, so the canonical bundle formula writes
(5) $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H).$
Let $\Sigma$ be a reduced snc divisor on $G$ and let $\overline{\Sigma}$ be its closure in $\mathbb{P}_{A}$. Then the canonical bundle formula for $(W,D_{W}+g^{*}\overline{\Sigma})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}+g^{*}\overline{\Sigma}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H+\overline{\Sigma})$. By inversion of adjunction (cf. [Fil20, Proposition 4.16]), $(W,D_{W}+g^{*}\overline{\Sigma})|_{g^{-1}G}$ is lc. Hence the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable by [Kol23, Corollary 4.55]. In particular, the restriction $g|_{g^{-1}G}$ is equidimensional, and so $W$ has no exceptional divisors over $G$.
Let $Q\subset D_{W}^{v}$ be a prime divisor. If $g(Q)\not\subset H$, then $g(Q)$ is a divisor on $\mathbb{P}_{A}$. But by the canonical bundle formula (5), we must have $H\geq g(Q)$, which is a contradiction. Hence $D_{W}^{v}=g^{-1}H$.
(5) By (3), it suffices to show that for any prime divisor $P\subset H$, there is a prime divisor $Q\not\subset E_{Z}$ dominating $P$. Assume the contrary, i.e., assume that there is a prime divisor $P\subset H$, and an effective exceptional over $\mathbb{P}_{A}$ divisor $E$ with $f(E)\subset P$ such that
(6) $f^{-1}P\subset E_{Z}+E.$
Case 1. $\dim\mathbb{P}_{A}-\dim A=1$. By [BFPT24, Theorem 1.5], the Albanese morphism $p_{X}:X\to A^{\prime}$ of $X$ is a fibration. Since $X$ has rational singularities, we have $\dim A^{\prime}=q(X)=q(Z)=\dim A$. Hence, the induced morphism $A\to A^{\prime}$ is an isomorphism. We identify $A^{\prime}$ with $A$. We have the following commutative diagram.
[…]
where $p$ is the natural projection.
We claim that
(7) $f^{-1}P\subset E_{Z}.$
Let $Q\subset E$ be a prime divisor. Then $Q\subset f^{-1}H\subset D_{Z}$. Since $\dim\mathbb{P}_{A}-\dim A=1$, $f(Q)$ is vertical over $A$, so $f(Q)$ is not an lc center of $(X,D_{X})$ by [BFPT24, Theorem 1.5]. Hence
$1<a(Q,X,D_{X})=-\operatorname{mult}_{Q}(D_{Z}-E_{Z})=-1+\operatorname{mult}_{Q}E_{Z}.$
Hence $Q\subset E_{Z}$ and the claim follows.
Let $L_{A}$ be an effective and sufficiently ample divisor on $A$. Since $P$ is ample over $A$, we have
$\dim\mathbb{P}_{A}$ $=\kappa(P+p^{*}L_{A})$
$=\kappa(f^{*}P+(p\circ f)^{*}L_{A})$
[…]
$=\kappa(E_{Z}+(p_{X}\circ\pi)^{*}L_{A})$
$=\kappa(p_{X}^{*}L_{A})\text{ ($E_{Z}$ is exceptional over $X$)}$
$=\dim A,$
which is a contradiction, so the inclusion (6) can not be true.
Case 2. $\dim\mathbb{P}_{A}-\dim A>1$. Since $P\subset H=\sum_{i}p_{i}^{*}H_{i}$, we have $P\subset p_{i}^{*}H_{i}$ for some $i$. By re-indexing $i$’s, we may assume that $P\subset p_{1}^{*}H_{1}$. Let
$\mathbb{P}^{\prime}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{2})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}),$
let $p_{i}^{\prime}:\mathbb{P}_{A}^{\prime}\to\mathbb{P}_{A}(O_{A}\oplus L_{i})$, $2\leq i\leq d$, $\alpha:\mathbb{P}_{A}\to\mathbb{P}_{A}^{\prime}$ be the natural projections and let $H^{\prime}=\sum_{i=2}^{d}{p^{\prime}}_{i}^{*}H_{i}$. Then
$(\alpha\circ f)^{-1}H^{\prime}\subset f^{-1}H\subset D_{Z}.$
By induction on $(\dim\mathbb{P}_{A}-\dim A)$, $(Z,D_{Z})$ has a good minimal model $(V,D_{V})$ over $\mathbb{P}_{A}^{\prime}$ with $K_{V}+D_{V}\sim_{\mathbb{Q}}0$. We have the following commutative diagram.
[…]
where $p^{\prime}$ is the natural projection and $p_{V}$ is the Albanese morphism of $V$ (by the same reasoning as in Case 1, we may identify the Albanese varieties of $Z$ and $V$).
We claim that
(8) $f^{-1}P\subset E_{Z}+(\alpha\circ f)^{-1}H^{\prime}.$
Let $Q\subset E$ be a prime divisor which is not contained in $E_{Z}$. Then $Q\subset f^{-1}H\subset D_{Z}$. Since $Z\dashrightarrow V$ contracts precisely the divisor $E_{Z}$, the strict transform $Q_{V}$ of $Q$ on $V$ is a divisor contained in $D_{V}$. Since $\dim\mathbb{P}_{A}-\dim\mathbb{P}_{A}^{\prime}=1$, and $Q$ is exceptional over $\mathbb{P}_{A}$, we have $Q$ is vertical over $\mathbb{P}_{A}^{\prime}$. Hence $Q_{V}$ is vertical over $\mathbb{P}_{A}^{\prime}$. By (4), we have $g(Q_{V})\subset H^{\prime}$. Hence $Q\subset(\alpha\circ f)^{-1}H^{\prime}$ and the claim follows.
Let $L_{A}$ be an effective and sufficiently ample divisor on $A$. Since $H^{\prime}$ and $\alpha^{*}H^{\prime}+P$ are ample over $A$, we have
$\dim\mathbb{P}_{A}$ $=\kappa(P+\alpha^{*}H^{\prime}+(p^{\prime}\circ\alpha)^{*}L_{A})$
$=\kappa(f^{*}P+(\alpha\circ f)^{*}H^{\prime}+(p^{\prime}\circ\alpha\circ f)^{*}L_{A})$
[…]
$=\dim\mathbb{P}_{A}^{\prime},$
which is a contradiction, so the inclusion (6) can not be true. ∎
Lemma 16.
Let $G$ be a quasi-abelian variety and let $P\subset G$ be a prime divisor, then $\overline{\kappa}(G\setminus P)>0$.
Proof.
The idea is from [Fuj24, Proof of Theorem 1.3, Step 2]. Let $A$ be the abelian part of $G$ and let $p:G\to A$ be the natural projection.
Case 1. $P$ is vertical over $A$. Then $P=p^{*}Q$ for some prime divisor $Q$ on $A$. Let $\mathbb{P}_{A}\to A$ be a $(\mathbb{P}_{1})^{d}$-bundle over $A$ such that $\mathbb{P}_{A}$ is a compactification of $G$, and $H=\mathbb{P}_{A}\setminus G$ is a snc divisor with $K_{\mathbb{P}_{A}}+H\sim 0$. Let $\overline{P}$ be the closure of $P$ in $\mathbb{P}_{A}$. Since any lc center of $(\mathbb{P}_{A},H)$ is horizontal over $A$, $\overline{P}$ contains no lc centers of $(\mathbb{P}_{A},H)$. Then $\overline{\kappa}(G\setminus P)=\kappa(K_{\mathbb{P}_{A}}+H+\overline{P})=\kappa(\overline{P})=\kappa(Q)>0$ by Lemma 17.
Case 2. $P$ is horizontal over $A$. Then there is a subgroup $\mathbb{C}^{*}\subset G$ such that $P$ dominates the quasi-abelian variety $G^{\prime}:=G/\mathbb{C}^{*}$. Let $F$ be a general fiber of $G\setminus P\to G^{\prime}$. Then $F$ is a curve with $\overline{\kappa}(F)>0$. Hence $\overline{\kappa}(G\setminus P)\geq\overline{\kappa}(F)+\overline{\kappa}(G^{\prime})=\overline{\kappa}(F)>0$ by Kawamata [Kaw78, Theorem 1] (see also Fujino [Fuj20, Chapter 5]).∎
Lemma 17.
Let $(Y,D_{Y})$ be a projective log smooth pair with $D_{Y}$ reduced. Assume that $\kappa(K_{Y}+D_{Y})\geq 0$. Let $\Sigma\subset Y$ be a reduced divisor which contains no lc centers of $(Y,D_{Y})$, then $\overline{\kappa}(Y\setminus D_{Y}+\Sigma)=\kappa(K_{Y}+D_{Y}+\Sigma)$.
Proof.
Let $\mu:Z\to Y$ be a log resolution of $(Y,D_{Y}+\Sigma)$. We may assume that $\mu$ is an isomorphism over $Y\setminus D_{Y}+\Sigma$. Let $\epsilon>0$ be a sufficiently small rational number. Then
$\overline{\kappa}(Y\setminus D_{Y}+\Sigma)$ $=\overline{\kappa}(Z\setminus\mu^{-1}(D_{Y}+\Sigma))$
$=\kappa(K_{Z}+\mu^{-1}(D_{Y}+\Sigma))$
$=\kappa(K_{Z}+\mu^{-1}_{*}D_{Y}+\mu^{-1}_{*}\Sigma+\operatorname{Ex}(\mu))$
$=\kappa(K_{Z}+\mu^{-1}_{*}D_{Y}+\epsilon\mu^{-1}_{*}\Sigma+\operatorname{Ex}(\mu))\text{ (by $\kappa(K_{Z}+\mu^{-1}_{*}D_{Y}+\operatorname{Ex}(\mu))\geq 0$)}$
$=\kappa(K_{Y}+D_{Y}+\epsilon\Sigma)\text{ (by $(Y,D_{Y}+\epsilon\Sigma)$ is lc)}.$
$=\kappa(K_{Y}+D_{Y}+\Sigma)\text{ (by $\kappa(K_{Y}+D_{Y})\geq 0$)}.$
Lemma 18.
Let $\alpha:Z\to V$ and $\beta:W\to V$ be fibrations between projective normal varieties, and let $Z\dashrightarrow W$ be a birational contraction over $V$. Let $E$ be an effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $Z$ which is exceptional over $W$, and let $L$ be a $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $V$. Then $\kappa(\alpha^{*}L+E)=\kappa(L)$.
Proof.
Let $u:X\to Z$ be a birational morphism resolving the indeterminacies of $Z\dashrightarrow W$ and let $v:X\to W$ be the induced morphism. Since $Z\dashrightarrow W$ is a birational contraction, $u^{*}E$ is exceptional over $W$. Hence
$\kappa(\alpha^{*}L+E)=\kappa(u^{*}\alpha^{*}L+u^{*}E)=\kappa(v^{*}\beta^{*}L+u^{*}E)=\kappa(\beta^{*}L)=\kappa(L).$
Theorem 2 is a special case of the following theorem.
Theorem 19.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $\alpha:X\setminus D\to G$ be the quasi-Albanese morphism. Then the conclusion in Theorem 2 holds.
Proof.
Keep the notations introduced in the proof of Theorem 3.
(1) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$. Then $\overline{P}\not\subset H$. Since $D_{W}^{v}=g^{-1}H$, there is a prime divisor $\overline{Q}_{W}\subset W$ such that $\overline{Q}_{W}\not\subset D_{W}$ and $\overline{Q}_{W}$ dominates $\overline{P}$. Let $\overline{Q}$ be the strict transform of $\overline{Q}_{W}$ on $X$. Then $\overline{Q}\not\subset D_{X}$ and $\overline{Q}$ dominates $P$.
(2) Let $Q\subset X\setminus D_{X}$ be a prime divisor and let $\overline{Q}$ be its closure in $X$. Then $\overline{Q}\not\subset D_{X}$ and its strict transform $\overline{Q}_{W}$ on $W$ has $\overline{Q}_{W}\not\subset D_{W}$, so $\overline{Q}_{W}\not\subset g^{-1}H$. Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable and in particular equidimensional, we have $\operatorname{codim}\alpha(Q)=\operatorname{codim}g(\overline{Q}_{W})=1$.
(3) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$. Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, we have $(W,D_{W}+g^{*}\overline{P})$ is lc near the generic point of $\overline{P}$, so $(Z,D_{Z}-E_{Z}+f^{*}\overline{P})|_{f^{-1}G}$ is sub-lc near the generic point of $P$. By
$(K_{Z}+D_{Z}-E_{Z}+f^{*}\overline{P})|_{Z\setminus\pi^{-1}D_{X}}=(\pi|_{Z\setminus\pi^{-1}D_{X}})^{*}(K_{X\setminus D_{X}}+\alpha^{*}P)$
and $Z\setminus\pi^{-1}D_{X}\subset Z\setminus f^{-1}H=f^{-1}{G}$, we have $(X\setminus D_{X},\alpha^{*}P)$ is lc near the generic point of $P$. ∎
4. A remark on orbifold fundamental groups
This section is inspired by the work [GLM23]. Let $(X,D_{X}=\sum_{i}(1-\frac{1}{m_{i}})D_{i})$ be an lc pair with standard coefficients, where each $D_{i}$ is a prime divisor and $m_{i}\in\mathbb{Z}_{>0}\cup\infty$. For each $i$, let $\gamma_{i}$ be a small loop around $D_{i}$. The orbifold fundamental group $\pi_{1}^{\operatorname{orb}}(X,D_{X})$ is defined as the quotient of $\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$ by the normal subgroup generated by $\gamma_{i}^{m_{i}}$’s. Note that if $D_{X}$ is reduced, then $\pi_{1}^{\operatorname{orb}}(X,D_{X})=\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$.
Conjecture 20.
Let $(X,D_{X})$ be an lc pair with standard coefficients. If $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, then $\pi_{1}^{\operatorname{orb}}(X,D)$ is virtually nilpotent.
Using the index-one cover and the Galois correspondence for orbifold fundamental groups (cf. [GLM23, Proposition A.11]), Conjecture 20 is reduced to the case $D_{X}$ is reduced. By the following Lemma, Conjecture 20 is further reduced to the $\mathbb{Q}$-factorial dlt case.
Lemma 21.
Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective lc pairs with $D_{X}$ and $D_{Y}$ reduced. If there is a birational contraction $p:X\dashrightarrow Y$ such that $D_{X}=p^{-1}_{*}D_{Y}+\operatorname{Ex}(p)$, then there is an isomorphism $\pi_{1}^{\operatorname{orb}}(X,D_{X})\cong\pi_{1}^{\operatorname{orb}}(Y,D_{Y})$.
Proof.
Let $Z\subset X$ and $Z^{\prime}\subset Y$ be closed subsets such that $X\setminus Z\cong Y\setminus Z^{\prime}$. Hence $X^{\operatorname{reg}}\setminus Z\cong Y^{\operatorname{reg}}\setminus Z^{\prime}$ and $X^{\operatorname{reg}}\setminus D_{X}\cup Z\cong Y^{\operatorname{reg}}\setminus D_{Y}\cup Z^{\prime}$. Since $p$ is a birational contraction, we have $\operatorname{codim}Z^{\prime}\geq 2$ and $D_{X}\cup Z=D_{X}\cup Z^{\prime\prime}$ for some $\operatorname{codim}\geq 2$ closed subset $Z^{\prime\prime}$. Hence $\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})\cong\pi_{1}(X^{\operatorname{reg}}\setminus D_{X}\cup Z^{\prime\prime})\cong\pi_{1}(X^{\operatorname{reg}}\setminus D_{X}\cup Z)\cong\pi_{1}(Y^{\operatorname{reg}}\setminus D_{Y}\cup Z^{\prime})\cong\pi_{1}(Y^{\operatorname{reg}}\setminus D_{Y})$. ∎
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
By a theorem due to Arapura and Nori [AN99, Corollary 4.10], to show that $\pi_{1}^{\text{orb}}(X,D_{X})=\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$ is virtually nilpotent, it suffices to show that it is polycyclic. Recall that a group $G$ is called polycyclic if there is a series of subgroups $1=G_{n}\triangleleft G_{n-1}\triangleleft\cdots\triangleleft G_{0}=G$ such that each $G_{i}/G_{i+1}$ is cyclic. The advantage of the class of polycyclic groups is that they are closed under passing to subgroups, quotients, and extensions (cf. [KM79, 4.4.3. Exercise]), which is crucial to inductive arguments.
Assume that $q(X,D_{X})>0$, and let $g:(W,D_{W})\to\mathbb{P}_{A}$ be the fibration given by Theorem 3. By Lemma 21, $(X,D_{X})$ and $(W,D_{W})$ have the same orbifold fundamental group. Let $W^{0}=g^{-1}G$ and let $F$ be a general fiber of $g$. Then $(F,{D_{W}}|_{F})$ is lc and $K_{F}+D_{W}|_{F}\sim_{\mathbb{Q}}0$. By the following exact sequence (cf. [GLM23, Lemma 2.18])
$\pi_{1}^{\operatorname{orb}}(F,D_{W}|_{F})\to\pi_{1}^{\operatorname{orb}}(W^{0},{D_{W}}|_{W^{0}})\to\pi_{1}(G)\to 1,$
we have $\pi_{1}^{\operatorname{orb}}(X,D_{X})\cong\pi_{1}^{\operatorname{orb}}(W,D_{W})\cong\pi_{1}^{\operatorname{orb}}(W^{0},{D_{W}}|_{W^{0}})$ is polycyclic if $\pi_{1}^{\operatorname{orb}}(F,D_{W}|_{F})$ is polycyclic. Hence, we have shown the following by induction on dimension.
Proposition 22.
To prove Conjecture 20, it suffices to prove $\mathbb{Q}$-factorial dlt pairs $(X,D_{X})$ with $D_{X}$ reduced and $q(X,D_{X})=0$ have polycyclic orbifold fundamental groups.